# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

### v3

@@ 1 Introduction
These MARL interaction paradigms cannot be fully captured by the conventional MARL learning paradigms (i.e., independent learning (IL) (albrecht2024multi, )[Ch. 9.3] and centralized training with decentralized execution (CTDE) (albrecht2024multi, )[Ch. 9.1.3]).
@@
While beneficial, applying global guidance to the entire multi-agent system in MARL is often impractical due to complexity, costs liang2024learning and safety concerns garg2024learning . Consider, for example, autonomous vehicles navigating a complex intersection or merging onto a highway. Ensuring safe and efficient passage requires intricate coordination. Also, providing simultaneous, specific feedback or instructions to all vehicles is infeasible, due to challenges such as complex safety validation and lack of universal communication protocols taeihagh2019governing . Yet, effective group coordination remains essential. An intriguing possibility arises if guidance is provided to a single targeted vehicle, e.g., instructing it to adjust its speed or yield at a critical moment. The resulting behaviour can serve as a pattern, enabling surrounding vehicles to coordinate. This scenario highlights a critical challenge where global guidance is essential but prohibitively difficult or costly, leading directly to the central question motivating our work:
@@
Can effective coordination still be achieved by assigning an additional desired outcome to a single agent, relying on its influence over the rest of the agents in a multi-agent system?
@@
Addressing the challenge of designing effective targeted intervention requires a principled framework to model and analyze how guiding a single agent impacts the overall multi-agent system. To this end, we propose leveraging multi-agent influence diagrams (MAIDs) koller2003multi , as a formal graphical language to encode complex strategic dependencies and information flow within the MARL learning process. Crucially, their bundled relevance graphs enable the analysis of a key property that we refer to as solvability. The identification of solvability guides better understanding and prediction of MARL interaction paradigms. To answer our research question, we utilize MAIDs to design the targeted intervention paradigm that intervenes on a single targeted agent, which reliably guides the entire multi-agent system towards a desired outcome. Formally, we frame this problem as selecting a desired Nash equilibrium among many possible equilibria harsanyi1988general , ensuring both primary task completion and satisfaction of an additional desired outcome.
@@
The main contributions of this paper are summarized as follows: (1) We introduce the concept of MARL interaction paradigms, using MAIDs to analyze and visualize both unguided self-organization and global guidance mechanisms in MARL. Then, we propose a new MARL interaction paradigm, referred to as the targeted intervention paradigm, where only a single targeted agent is intervened on. (2) The solvability of existing MARL learning paradigms under various MARL interaction paradigms can be analyzed, with the help of visualizing the corresponding relevance graphs. (3) Since MAIDs can be viewed as causal diagrams, we draw on the principle of causal inference to implement the targeted intervention paradigm, which we refer to as the Pre-Strategy Intervention (PSI). It is provable that the PSI can reach a composite desired outcome consisting of the primary task goal and an additional desired outcome. (4) The PSI is implemented as a pre-policy module that can be integrated into generic MARL algorithms.
@@ 2 Multi-Agent Influence Diagrams (MAIDs)
We now review multi-agent influence diagrams (MAIDs) (koller2003multi, ), which is an augmentation of the Bayesian network to describe multi-agent decision making to maximize their utility. In the rest of the paper, the terms variable and node will be used interchangeably. An MAID is usually described as a tuple $\mathcal{M}=(\mathcal{I},\mathcal{X},\mathcal{D},\mathcal{U},\mathcal{G},Pr)$, where $\mathcal{I}$ is a set of agents. $\mathcal{X}$ is a set of chance variables indicating decisions of nature. Each chance variable $X\in\mathcal{X}$ is associated with a set of parents $Pa(X)\subset\mathcal{X}\cup\mathcal{D}$. The $\mathcal{D}:=\bigcup_{i\in\mathcal{I}}\mathcal{D}_{i}$ is a set of all agents’ decision variables, where $\mathcal{D}_{i}$ is the set of agent $i$’s decision variables. For a decision variable $D\in\mathcal{D}_{i}$, $Pa(D)$ is the set of variables whose values are informed to agent $i$ when it selects a value of $D$. $\mathcal{U}:=\bigcup_{i\in\mathcal{I}}\mathcal{U}_{i}$ is a set of utility variables, where $\mathcal{U}_{i}$ is agent $i$’s utility variable as its utility function. Note that utility variables cannot be parents of other variables. A directed acyclic graph $\mathcal{G}$ with variables $\mathcal{V}=\mathcal{X}\cup\mathcal{D}\cup\mathcal{U}$ is formed. $Pr$ is a conditional probability distribution (CPD) defined over chance variables $X$ such as $Pr(X|Pa(X))$, and utility variables $U\in\mathcal{U}$ such as $Pr(U|\mathbf{pa})$, for each $\mathbf{pa}\in dom(Pa(U))$. Note that $Pr(U|Pa(U))$ is a Dirac function (i.e. $U$ is a deterministic function). In other words, for each instantiation $\mathbf{pa}\in dom(Pa(U))$, there is a value of $U$ that is assigned probability 1, and probability of other values is 0. To simplify the notation, $U(\mathbf{pa})$ denotes as the value of $U$ that has probability 1 when $Pa(U)=\mathbf{pa}$. The total utility that an agent $i$ obtained from an instantiation of $\mathcal{V}$ is the sum of the values of $\mathcal{U}_{i}$, i.e. $\sum_{U\in\mathcal{U}_{i}}U(\mathbf{pa})$ where $\mathbf{pa}\in dom(Pa(U))$.
@@
Let $\mathcal{E}$ be a subset of $\mathcal{D}_{i}$, and let $\sigma$ be a strategy profile. $\sigma_{\mathcal{E}}^{*}$ is optimal for the strategy profile $\sigma$ if, in the induced MAID $\mathcal{M}[\sigma_{-\mathcal{E}}]$, where the only remaining decisions are those in $\mathcal{E}$, the strategy $\sigma_{\mathcal{E}}^{*}$ is optimal, for all strategies $\sigma_{\mathcal{E}}^{\prime}$, such that
$\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{*}))\geq\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{\prime})).$
@@
A node $D^{\prime}$ is s-reachable from a node $D$ in a MAID $\mathcal{M}$ if there is some utility node $U\in\mathcal{U}_{D}$ such that if a new parent $\hat{D}^{\prime}$ were added to $D^{\prime}$, there would be an active path (see Example C.1 and Definition C.1 in Appendix C.2) in $\mathcal{M}$ from $\hat{D}^{\prime}$ to $U$ given $Pa(D)\cup\{D\}$, where a path is active in a MAID if it is active in the same graph, viewed as a Bayesian network (i.e. it is not blocked as per the standard d-separation rules). The relevance graph for a MAID $\mathcal{M}$ is a directed graph whose nodes are the decision nodes of $\mathcal{M}$, and which contains an edge $D^{\prime}\rightarrow D$ if and only if $D^{\prime}$ is s-reachable from $D$.
Relevance Graph. A relevance graph as shown in Definition 2.4 defines a directed graph induced from an MAID, describing the binary relation between two decision variables. If there exists an edge $D^{\prime}\rightarrow D$, it implies that the decision variable $D$ is strategically relies on another decision variable $D^{\prime}$. In other words, the decision rules for $D^{\prime}$ is required to evaluate the decision rules for $D$. If there exist both $D^{\prime}\rightarrow D$ and $D\rightarrow D^{\prime}$, then the relevance graph is cyclic. Furthermore, if $D$ and $D^{\prime}$ belong to two agents respectively, their payoffs depend on the decisions at both $D$ and $D^{\prime}$. Under this situation, the optimality of one agent’s decision rule is coupled with another agent’s decision rule.
@@
One way to address the above issue is making these two agents’ decision rules matched (koller2003multi, ), e.g. choosing both agents’ decision rules together.
@@ 3.1 Formalizing MARL Interaction Paradigms
By the definition of MAIDs in Section 2, we model MARL interaction paradigms based on agent variables ($\mathcal{I}$), decision variables ($\mathcal{D}$), chance variables ($\mathcal{X}$) representing states and information, and utility variables ($\mathcal{U}$) representing objectives. Utility variables capture the desired outcomes that agents pursue, whereas the special chance variables $\mathcal{Z}\subset\mathcal{X}$ describe signals (e.g. behaviours, communications, etc.) to guide agents for achieving their desired outcomes. The formal definition is delineated in Definition 3.1.
@@
(1) Self-Organization. This paradigm refers to the MARL setting where agents coordinate solely via their direct observations of the environment, without any external guidance or structural intervention to steer them towards their desired outcomes. Each agent must independently elicit a signal (a pattern of behaviours) based on local information to guide the other agent for achieving its desired outcome. For clarity of illustration, we focus on analyzing guidance to a single agent. We visualize this implicit process as a path from decision variables $D_{a}$ to the signal $Z$, and then to the utility nodes $U_{h}$, illustrated by MAID structures in Figure 1(a) and 2(a).
(2) Global Intervention. This paradigm features a external, centralized coordination providing simultaneous explicit guidance signals to all agents. In the MAID, as shown in Figure 1(b) and 2(b), this is represented as information links from $Z$ to agents’ decision variables, and $Z$ usually influence their utility nodes (e.g., $U_{a}$ and $U_{h}$) directly, shaping their desired outcomes.
(3) Targeted Intervention. This paradigm involves applying an external guidance signal to a targeted agent, as shown in Figure 1(c) and 2(c). Specifically, $D_{pre}$ receives a guidance signal $Z$ and outputs filtered information. This intervention modifies the behaviours of the targeted agent, which in turn indirectly influences other agents. As a result, agents coordinate to achieve their desired outcomes provided by the guidance signal $Z$. While the concept of targeted intervention paradigm could potentially apply to multiple targeted agents with multiple desired outcomes, this paper specifically focuses on the case of a single targeted agent.
@@ 3.2 Interpreting MARL Interaction Paradigms
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi . of MARL learning algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the self-organization paradigm are cyclic, implying computational or theoretical intricacies in finding solutions. In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependencies colas2022autotelicagentsintrinsicallymotivated . While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on the targeted intervention paradigm provides a distinct advantage: it achieves solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention paradigm.
@@ 3.3 Pre-Strategy Intervention: An Implementation of Targeted Intervention Paradigm
This section introduces the pre-strategy intervention—a principled design of the targeted intervention paradigm. Through the pre-strategy intervention, we aim to guide the multi-agent system towards a preferred Nash equilibrium (NE) that satisfies both the primary task goal and an additional desired outcome. Our approach first defines the pre-strategy intervention, then details how its influence on achieving this NE is quantified and optimized via learning a pre-policy to generate pre-strategies intervening on a single targeted agent.
@@ 3.3.1 Definition of Pre-Strategy Intervention
To regulate the intervened agent to maximize the desired outcome $U$, we propose to add a pre-decision variable $D_{pre}$ as a new parent to the decision variable $D_{h}$ of a selected agent $h\in\mathcal{I}$. In analogy to decision variables in MAIDs, we need to assign a strategy $\sigma_{pre}$ which we refer to as pre-strategy, and this operation is named pre-strategy intervention. This definition respects the convention of stochastic intervention in causal Bayesian networks (pearl2009causality, )[Ch. 4]. A pre-strategy is determined by a pre-policy denoted by $\delta_{pre}$, which processes the agent’s information and the guidance signal, as detailed in Definition 3.2.
@@ 3.3.2 Causal Effect of Pre-Strategy Intervention
To design the pre-strategy intervention, we first define how its causal effect is measured. The composite desired outcome is captured by a total utility variable $U_{tot}:=U_{tot}^{sec}+U_{tot}^{task}$, comprising the primary task goal $U_{tot}^{task}$ and an additional desired outcome (a secondary goal) $U_{tot}^{sec}$. We formalize the causal effect of a pre-strategy intervention in Definition 3.3. The causal effect quantifies the total probabilities of $U_{tot}=u^{*}$, where $u^{*}$ is the optimal utility value of the desired composite outcome, under a pre-strategy intervention. In many cases, it is difficult to find a pre-strategy intervention that induces the optimal NE. Instead, we seek a pre-strategy that induces a probability distribution over a reduced set of NEs covering the optimal NE, denoted as $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$. Proposition 3.4 proves pre-strategy intervention maximizing this causal effect is guaranteed to exist.
@@
A pre-strategy intervention is applied to a targeted agent’s strategy profiles, which influences the total utility $U_{tot}=u^{*}$ associated with a Nash equilibrium (NE) $\hat{\sigma}$ after the pre-strategy intervention. The set of all possible NEs is denoted by $\boldsymbol{\hat{\sigma}}$ (with no interventions). Given that $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$ denotes a reduced set of NEs induced by a pre-strategy intervention, covering the optimal NE (aligned to the additional desired outcome), the causal effect measuring this pre-strategy intervention is defined as follows:
@@
To practically evaluate possible strategy profiles determined by the pre-strategies generated by a pre-policy, we reshape $P_{\mathcal{I}}$ as the following expression:
@@
Note that $P_{\mathcal{U}}$ is a constant with respect to the intervention, so it can be ignored during optimization. In Eq. (3), $P_{\hat{\sigma}}(\hat{\sigma}\mid\operatorname{do}(\sigma_{pre}))$ is a probability distribution whose support includes all possible strategy profiles, which is feasible to implement and induce any possible $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$. Thus, we only need to maximize Eq. (3) in implementation.

### v4

@@ 1 Introduction
These MARL interaction paradigms go beyond the scope of the conventional MARL learning paradigms (e.g., centralized training and decentralized execution (CTDE) (albrecht2024multi)[Ch. 9.1.3]).
@@
While beneficial, applying global guidance to the entire multi-agent system in MARL is often impractical due to complexity, costs liang2024learning and safety concerns garg2024learning. Consider, for example, autonomous vehicles navigating a complex intersection or merging onto a highway. Ensuring safe and efficient passage requires intricate coordination. Also, providing simultaneous, specific feedback or instructions to all vehicles is infeasible, due to challenges such as complex safety validation and lack of universal communication protocols taeihagh2019governing. Yet, effective group coordination remains essential. An intriguing possibility arises if an additional desired outcome as a guidance is applied to a single targeted vehicle, e.g., instructing it to adjust its speed or yield at a critical moment. The resulting behaviour can serve as a pattern, enabling surrounding vehicles to coordinate. This scenario highlights a critical challenge where global guidance is essential but prohibitively difficult or costly, leading directly to the central question motivating our work:
@@
Can effective coordination be achieved when assigning an additional desired outcome to a single targeted agent, relying on its influence over the rest of the agents in a multi-agent system?
@@
Addressing the challenge of designing effective targeted intervention requires a principled framework to model and analyze how guiding a single agent impacts the overall multi-agent system. To this end, we propose leveraging multi-agent influence diagrams (MAIDs) koller2003multi, as a formal graphical language to encode complex strategic dependencies and information flow within the MARL learning process. Crucially, their bundled relevance graphs enable the analysis of a key property that we refer to as solvability. The identification of solvability guides better understanding and prediction of MARL interaction paradigms. To answer our research question, we utilize MAIDs to design the targeted intervention paradigm that intervenes on a single targeted agent, which reliably guides the entire multi-agent system towards a desired outcome. Formally, we frame this problem as selecting a preferred Nash equilibrium (NE) from many possible equilibria harsanyi1988general, where the preferred NE is precisely the one that ensures effective coordination with satisfying an additional desired outcome.
@@
The main contributions of this paper are summarized as follows: (1) We introduce the concept of MARL interaction paradigms, using MAIDs to analyze and visualize both unguided self-organization and global guidance mechanisms in MARL. Then, we propose a new MARL interaction paradigm, referred to as the targeted intervention paradigm, where only a single targeted agent is intervened on. (2) The solvability of existing MARL learning paradigms under various MARL interaction paradigms can be analyzed, with the help of visualizing the corresponding relevance graphs. (3) Since MAIDs can be viewed as causal diagrams, we draw on the principle of causal inference to implement the targeted intervention paradigm, which we refer to as the Pre-Strategy Intervention (PSI). It is provable that the PSI can reach a composite desired outcome consisting of the primary task goal for coordination and an additional desired outcome. (4) The PSI is implemented as a pre-policy module that can be integrated into generic MARL algorithms.
@@ 2 Multi-Agent Influence Diagrams (MAIDs)
We now review multi-agent influence diagrams (MAIDs) (koller2003multi), which is an augmentation of the Bayesian network to describe multi-agent decision making to maximize their utility. In the rest of the paper, the terms variable and node will be used interchangeably. An MAID is usually described as a tuple $\mathcal{M}=(\mathcal{I},\mathcal{X},\mathcal{D},\mathcal{U},\mathcal{G},Pr)$, where $\mathcal{I}$ is a set of agents. $\mathcal{X}$ is a set of chance variables indicating decisions of nature. Each chance variable $X\in\mathcal{X}$ is associated with a set of parents $Pa(X)\subset\mathcal{X}\cup\mathcal{D}$. The $\mathcal{D}:=\bigcup_{i\in\mathcal{I}}\mathcal{D}_{i}$ is a set of all agents’ decision variables, where $\mathcal{D}_{i}$ is the set of agent $i$’s decision variables. For a decision variable $D\in\mathcal{D}_{i}$, $Pa(D)$ is the set of variables whose values are informed to agent $i$ when it selects a value of $D$. $\mathcal{U}:=\bigcup_{i\in\mathcal{I}}\mathcal{U}_{i}$ is a set of utility variables, where $\mathcal{U}_{i}$ is agent $i$’s utility variable as its utility function. Note that utility variables cannot be parents of other variables. A directed acyclic graph $\mathcal{G}$ with variables $\mathcal{V}=\mathcal{X}\cup\mathcal{D}\cup\mathcal{U}$ is formed. $Pr$ is a conditional probability distribution (CPD) defined over chance variables $X$ such as $Pr(X|Pa(X))$, and utility variables $U\in\mathcal{U}$ such as $Pr(U|\mathbf{pa})$, for each $\mathbf{pa}\in dom(Pa(U))$. Note that $Pr(U|Pa(U))$ is a Dirac function (i.e. $U$ is a deterministic function). In other words, for each instantiation $\mathbf{pa}\in dom(Pa(U))$, there is a value of $U$ that is assigned probability 1, and probability of other values is 0. To simplify the notation, $U(\mathbf{pa})$ denotes as the value of $U$ that has probability 1 when $Pa(U)=\mathbf{pa}$. The total utility that an agent $i$ obtained from an instantiation of $\mathcal{V}$ is the sum of the values of $\mathcal{U}_{i}$, i.e. $\sum_{U\in\mathcal{U}_{i}}U(\mathbf{pa})$ where $\mathbf{pa}\in dom(Pa(U))$.
@@
Let $\mathcal{E}$ be a subset of $\mathcal{D}_{i}$, and let $\sigma$ be a strategy profile. $\sigma_{\mathcal{E}}^{*}$ is optimal for the strategy profile $\sigma$ if, in the induced MAID $\mathcal{M}[\sigma_{-\mathcal{E}}]$, where the only remaining decisions are those in $\mathcal{E}$, the strategy $\sigma_{\mathcal{E}}^{*}$ is optimal, for all strategies $\sigma_{\mathcal{E}}^{\prime}$, such that $\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{*}))\geq\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{\prime}))$.
@@
A node $D^{\prime}$ is s-reachable from a node $D$ in a MAID $\mathcal{M}$ if there is some utility node $U\in\mathcal{U}_{D}$ such that if a new parent $\hat{D}^{\prime}$ were added to $D^{\prime}$, there would be an active path (Example C.1 and Definition C.1 in Appendix C.2) in $\mathcal{M}$ from $\hat{D}^{\prime}$ to $U$ given $Pa(D)\cup\{D\}$, where a path is active in a MAID if it is active in the same graph, viewed as a Bayesian network (i.e. it is not blocked as per the standard d-separation rules). The relevance graph for a MAID $\mathcal{M}$ is a directed graph whose nodes are the decision nodes of $\mathcal{M}$, and which contains an edge $D^{\prime}\rightarrow D$ if and only if $D^{\prime}$ is s-reachable from $D$.
Relevance Graph. A relevance graph as shown in Definition 2.4 defines a directed graph induced from an MAID, describing the binary relation between two decision variables. If there exists an edge $D^{\prime}\rightarrow D$, it implies that the decision variable $D$ is strategically relies on another decision variable $D^{\prime}$. In other words, the decision rules for $D^{\prime}$ is required to evaluate the decision rules for $D$. If there exist both $D^{\prime}\rightarrow D$ and $D\rightarrow D^{\prime}$, then the relevance graph is cyclic. Furthermore, if $D$ and $D^{\prime}$ belong to two agents respectively, their payoffs depend on the decisions at both $D$ and $D^{\prime}$. In this case, the optimality of one agent’s decision rule is intertwined with another agent’s decision rule (koller2003multi). Therefore, agents seeking decision rules individually to reach an NE, analogous to independent learning (albrecht2024multi)[Ch. 9.3], is unsolvable, because mismatch between agents’ decision rules could happen.
@@
One way to make cyclic relevance graphs solvable is making agents’ decision rules matched via seeking agents’ decision rules collectively (koller2003multi), analogous to CTDE (albrecht2024multi)[Ch. 9.1.3].
@@ 3.1 Formalizing MARL Interaction Paradigms
By the definition of MAIDs in Section 2, we model MARL interaction paradigms based on agent variables ($\mathcal{I}$), decision variables ($\mathcal{D}$), chance variables ($\mathcal{X}$) representing states and information, and utility variables ($\mathcal{U}$) representing objectives. Utility variables capture the desired outcomes that agents pursue, whereas the special chance variables $\mathcal{Z}\subset\mathcal{X}$ describe signals to guide agents for achieving their desired outcomes. The formal definition is delineated in Definition 3.1.
@@
(1) Self-Organization. This paradigm refers to the MARL setting where agents coordinate solely via their direct observations of the environment, without any external mechanism to steer them towards their desired outcomes. Each agent must independently elicit a signal (a pattern of behaviours) based on local information to influence the other agents towards its desired outcome. For clarity of illustration, we focus on analyzing guidance to a single agent. We visualize this implicit process as a path from decision variables $D_{a}$ to the signal $Z$, and then to the utility nodes $U_{h}$, illustrated by MAID structures in Figure 1(a) and 2(a).
(2) Global Intervention. This paradigm features an external, centralized coordination providing simultaneous explicit guidance signals to all agents. In the MAID, as shown in Figure 1(b) and 2(b), this is represented as information links from $Z$ to agents’ decision variables, and $Z$ usually influence their utility nodes (e.g., $U_{a}$ and $U_{h}$) directly, shaping their desired outcomes.
(3) Targeted Intervention. This paradigm involves applying an external guidance signal to a targeted agent, as shown in Figure 1(c) and 2(c). For conciseness, we illustrate this as one of its implementation, referred to as pre-strategy intervention (Section 3.3), where $D_{pre}$ receives a guidance signal $Z$ and outputs filtered information. This intervention modifies the behaviours of the targeted agent, which in turn indirectly influences other agents. As a result, agents coordinate to achieve their desired outcomes provided by the guidance signal $Z$. While the concept of targeted intervention paradigm could potentially apply to multiple targeted agents with multiple desired outcomes, this paper specifically focuses on the case of a single targeted agent.
@@ 3.2 Interpreting MARL Interaction Paradigms
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi. of MARL algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the self-organization paradigm are cyclic, implying computational or theoretical intricacies in finding solutions. In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependency colas2022autotelicagentsintrinsicallymotivated. While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on the targeted intervention paradigm provides a distinct advantage: it can facilitate solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention paradigm.
@@ 3.3 Pre-Strategy Intervention: An Implementation of Targeted Intervention Paradigm
This section introduces the pre-strategy intervention—a principled design of the targeted intervention paradigm. By applying a pre-strategy intervention, we aim to guide the multi-agent system to a preferred Nash equilibrium (NE), selecting one with an additional desired outcome from the multiple equilibria that satisfy the primary task goal. Our approach first defines the pre-strategy intervention, then details how its influence on reaching the preferred NE is quantified and optimized via learning a pre-policy to generate pre-strategies intervening on a single, targeted agent.
@@ 3.3.1 Definition of Pre-Strategy Intervention
To regulate the intervened agent to maximize the desired outcome $U$ that captures the preferred NE, we propose to add a pre-decision variable $D_{pre}$ as a new parent to the decision variable $D_{h}$ of a selected agent $h\in\mathcal{I}$. In analogy to decision variables in MAIDs, we need to assign a strategy $\sigma_{pre}$ which we refer to as pre-strategy, and this operation is named pre-strategy intervention. This definition respects the convention of stochastic intervention in causal Bayesian networks (pearl2009causality)[Ch. 4]. A pre-strategy is determined by a pre-policy denoted by $\delta_{pre}$, which processes the agent’s information and the guidance signal, as detailed in Definition 3.2.
@@ 3.3.2 Causal Effect of Pre-Strategy Intervention
The preferred NE is identified by maximizing a composite desired outcome, formally defined by a total utility variable $U_{tot}:=U_{task}+U_{sec}$, comprising the utility $U_{task}$ of the primary task goal that all agents coordinate to achieve and the utility $U_{sec}$ of an additional desired outcome (a secondary goal) only assigned to a single, targeted agent. [^3] The additional desired outcome is constrained to be non-conflicting with the primary task goal. This preferred NE is one of multiple equilibria that all yield the identical, maximum value of the primary task utility $U_{task}$. We formalize the causal effect of a pre-strategy intervention in Definition 3.3. The causal effect quantifies the total probabilities of $U_{tot}=u^{*}$, where $u^{*}$ is the utility value of the desired composite outcome, under a pre-strategy intervention for the additional desired outcome. In many cases, it is difficult to find a pre-strategy intervention that induces the preferred NE. Instead, we seek a pre-strategy that induces a probability distribution over a reduced set of NEs covering the preferred NE, denoted as $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$. Proposition 3.4 proves pre-strategy intervention maximizing this causal effect is guaranteed to exist.
@@
A pre-strategy intervention is applied to a targeted agent’s strategy profiles, which influences the total utility $U_{tot}=u^{*}$ associated with a Nash equilibrium (NE) $\hat{\sigma}$ after the pre-strategy intervention. The set of all possible NEs is denoted by $\boldsymbol{\hat{\sigma}}$ (with no intervention). Given that $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$ denotes a reduced set of NEs induced by a pre-strategy intervention, covering the preferred NE (aligned to the additional desired outcome), the causal effect measuring this pre-strategy intervention is defined as follows:
@@
To practically evaluate possible strategy profiles of the pre-strategies generated by a pre-policy, we reshape $P_{\mathcal{I}}$ as the following expression:
@@
Note that $P_{\mathcal{U}}$ is a constant with respect to the intervention, so it can be ignored during optimization. In Eq. (3), $P_{\hat{\sigma}}(\hat{\sigma}\mid\operatorname{do}(\sigma_{pre}))$ is a probability distribution whose support includes all possible strategy profiles, which is feasible to implement and induce any possible $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$.